A curling rock with a mass of 20.0 kg slides down a sheet of ice at a speed of 6.00 m/s. What is the kinetic energy of the curling rock?
1
Expert's answer
2020-12-21T11:02:57-0500

By definition, the kinetic energy is


"K = \dfrac{mv^2}{2}"

where "m = 20kg" is the mass of the rock, and "v = 6m\/s" is its speed. Thus, obtain:


"K = \dfrac{20\cdot 6^2}{2} = 360J"

Answer. 360 J.